HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

THE BENDING OF C_{3}: EXPERIMENTALLY PROBING THE l-TYPE DOUBLING AND RESONANCE

\underline{M.-A. MARTIN-DRUMEL} , O. PIRALI, M. VERVLOET, Université Paris-Saclay, CNRS, Institut des Sciences Moléculaires d’Orsay, France; D. W. TOKARYK, Department of Physics, University of New Brunswick, Fredericton, NB, Canada; K. D. DONEY, H. LINNARTZ, Leiden Observatory, Laboratory for Astrophysics, Universiteit Leiden, Leiden, Netherlands; Q. ZHANG,Y. CHEN,D. ZHAO, Leiden Observatory, Laboratory for Astrophysics, Universiteit Leiden,

Leiden, Netherlands

C_{3}, a pure carbon chain molecule that has been identified in different astronomical environments, is considered a good probe of kinetic temperatures through observation of transitions involving its low-lying bending mode (ν_{2}) in its ground electronic state. With the aim to investigate this bending mode with multiple quanta of excitation, we have undertaken high resolution optical and mid-infrared investigations on the species produced in discharge experiments.

We report here the most complete analysis of the X̃ and Ã states of C_{3} (literature and present study) using a single PGOPHER file. New experimental measurements result in 36 rovibronic Ã ^{1}Π_{u} − X̃ ^{1}Σ_{g}^{+} bands (originating from X̃(0v_{2}0), v_{2} = 0 − 5, levels) recorded by laser induced fluorescence spectroscopy at the University of Science and Technology of China and the rovibrational ν_{3} band (revealing hot bands involving up to 5 quanta of excitation in v_{2}) recorded by Fourier-transform infrared spectroscopy using a globar source on the AILES beamline of the SOLEIL synchrotron facility.

The combined fit allows for the accurate determination of the rotational parameters and absolute energy levels of C_{3}, in particular for states involving the bending mode. The spectroscopic information derived from this work enables new interstellar searches for C_{3}, not only in the infrared and optical regions investigated here but also notably in the ν_{2} band region (around 63 cm^{−1}) where vibrational satellites can now be accurately predicted. This makes C_{3} a universal diagnostic tool to study very different astronomical environments, from dark and dense to translucent clouds.

p-number: p107  Submitted on Tue, 13 Jun 2023 16:58:23 +0200